# Measuring Ransomware Lateral Movement Susceptibility via Privilege-Weighted Adjacency Matrix Exponentiation Satyam Tyagi1 Ganesh Murugesan1

### v1

@@ 4.1 Lateral-Movement Susceptibility (LMSK)
Concept: This metric quantifies the overall ease of movement within the network. It answers the question: ”If we randomly select two different assets, what is the average probability that a successful lateral movement event can occur between them within $K$ hops?”
@@ 4.2 Blast-Radius Estimate (BREK)
Concept: This metric quantifies the expected scope of compromise from an attack. It answers the business-critical question: ”Assuming an attacker can start at any asset, what percentage of our entire network do we expect to be compromised?”
@@
$\text{BRE}_{K}=\left(\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}(P_{K})_{ij}\right)\times 100\%$ (4)
@@ 5.  Formal Properties of the Model
Overview. We prove monotonicity, boundedness, and convergence; derive a $K$-hop series upper bound and a single-path product lower bound; recover boolean reachability in the unweighted limit; characterize the limit as the least fixed point $P^{*}=F(P)=P\oplus(P\otimes W)$; and show all results persist under service-class lifting.
@@ 5.1 Monotonicity
This is guaranteed by the definition of the probabilistic union operator ($\lor_{p}$). This property ensures the model behaves logically, as the total probability of compromise only ever increases or remains constant as more paths are considered.
@@ 5.2 Boundedness
The values in the compromise probability matrix are always bounded within the interval $[0,1]$. The pivot potentials $\pi(s)$ are defined in $[0,1]$, and both the probabilistic composition ($\otimes$) and union ($\lor_{p}$) operators are closed over this interval. This property guarantees the stability of the model, ensuring that the outputs are always valid probabilities.
5.3 Convergence
The sequence of matrices $P_{1},P_{2},\dots$ is guaranteed to converge to a fixed-point matrix $P^{*}$, where further iterations no longer change the result. This is a direct consequence of the Monotonicity and Boundedness properties (a monotonically increasing sequence that is bounded from above must converge).
Unlike unweighted reachability, which converges in at most n1n-1 steps, the probabilistic model may require more iterations for the probability values to become fully stable. This is because paths longer than n1n-1 can still contribute a small but non-zero probability to the total, incrementally refining the final result. In practice, the algorithm is governed by a stopping rule: iterate until the change between successive matrices falls below a predefined tolerance threshold, $\varepsilon$, or a maximum number of hops $K$ is reached. Formally, the iteration stops when $\|P_{k+1}-P_{k}\|_{\infty}<\varepsilon$ or $k=K$.
By continuity of $\oplus$ and $\otimes$ on $[0,1]$, the limit $P^{*}$ satisfies $P^{*}=F(P^{*})$ with $F(P)=P\oplus(P\otimes W)$; thus $P^{*}$ is the least fixed point above $P_{1}$.
@@ 5.4 Refinement of Binary Reachability
Let $S_{K}$ be the final reachability matrix from the unweighted (Boolean) method, and $P_{K}$ be the final matrix from our probabilistic method. The following soundness relationship holds:
$(P_{K})_{ij}>0\;\Rightarrow\;(S_{K})_{ij}=1$
@@
5.5 Connection to Algebraic Path Problems
Our construction follows the algebraic-path template over $[0,1]$ using the probabilistic-sum (noisy-OR) as the “addition” $\oplus$ and ordinary multiplication as the “multiplication” $\otimes$:
xy= $x\oplus y\;=\;1-(1-x)(1-y),\qquad x\otimes y\;$=\;xy.
Matrix composition then uses these operations in the usual way:
$(X\otimes Y)_{ij}\;=\;\bigoplus_{m=1}^{n}X_{im}\otimes Y_{mj}\;=\;1-\prod_{m=1}^{n}\bigl{(}1-X_{im}Y_{mj}\bigr{)},$
which is exactly our probabilistic composition operator. (Here $\oplus$ corresponds to the $\lor_{p}$ operator used elsewhere.)
Remark (semiring-like, non-distributive).
The pair $(\oplus,\otimes)$ is associative, commutative, and monotone with identities 0 (for $\oplus$) and $1$ (for $\otimes$), but it is not distributive. For all $a,b,c\in[0,1]$,
$(a\otimes b)\,\oplus\,(a\otimes c)\ \geq\ a\otimes(b\oplus c),\qquad(a\otimes c)\,\oplus\,(b\otimes c)\ \geq\ (a\oplus b)\otimes c,$
with equality iff, respectively, $a\in\{0,1\}$ or $bc=0$, and $c\in\{0,1\}$ or $ab=0$. We do not rely on distributivity in our results.
Fixed-point view.
Let $F(P)=P\oplus(P\otimes W)$. Since $F$ is monotone on the complete lattice $0,1]^{n\times n}$ and bounded in $[0,1]$, iterating
$P_{k+1}\;=\;P_{k}\oplus(P_{k}\otimes W)$
computes the path-closure (least fixed point) up to either a prescribed hop limit $K$ or an $\varepsilon$-tolerance, grounding our iteration in standard algebraic-path theory.
5.6 K-hop Series Bound for Lateral Propagation
Define, for $K\geq 1$, the path-sum matrix
$U_{K}\;=\;A\;+\;AW\;+\;AW^{2}\;+\;\cdots\;+\;AW^{K-1},$
where products are the usual matrix products. The term $AW^{t}$ aggregates all paths of length t+1t{+}1 (one landing hop in $A$ followed by $t$ pivot steps in $W$).
Proposition 1 (K-hop series upper bound).
For all $K\geq 1$ and all $i\neq j$,
$(P_{K})_{ij}\;\leq\;\min\!\bigl{(}1,\ (U_{K})_{ij}\bigr{)},\qquad\text{and}\qquad(P_{K})_{ii}=1.$
Proof. Fix $i\neq j$. Let $\mathcal{P}_{\leq K}(i\!\to\!j)$ be the set of all directed paths from $i$ to $j$ of length at most $K$. By construction, $(P_{K})_{ij}$ is the probability of the union of the path events $\{E_{p}\}_{p\in\mathcal{P}_{\leq K}}$. By Boole’s union inequality, $\Pr(\bigcup_{p}E_{p})\leq\sum_{p}\Pr(E_{p})$. For a path of length $L\geq 1$ with vertices $i\to m_{1}\to\cdots\to m_{L-1}\to j$, the model assigns probability $A_{im_{1}}\prod_{r=1}^{L-1}W_{m_{r}m_{r+1}}$, so summing over all such paths of fixed length $L$ yields $(AW^{L-1})_{ij}$. Summing over $L=1,\dots,K$ gives $(U_{K})_{ij}$. Capping at $1$ preserves validity. The diagonal equality $(P_{K})_{ii}=1$ follows from the initialization $P_{1}=I\lor_{p}A$. $\square$
Corollary 1 (infinite-hops bound).
If the spectral radius $\rho(W)<1$, then the Neumann series converges and
$U_{\infty}\;:=\;\sum_{t=0}^{\infty}AW^{t}\;=\;A(I-W)^{-1},\qquad\text{hence}\qquad(P_{\infty})_{ij}\;\leq\;\min\!\bigl{(}1,\ (U_{\infty})_{ij}\bigr{)}.$
Proof. Since $\rho(W)<1$, $\sum_{t\geq 0}W^{t}=(I-W)^{-1}$ converges. Apply Proposition 1 with $K\to\infty$ and linearity of matrix multiplication. $\square$
@@ 5.6 K-hop Series Bound for Lateral Propagation
Fix $K\geq L\geq 1$. Suppose there exists a directed path $p:\ i\!=\!m_{0}\to m_{1}\to\cdots\to m_{L-1}\to m_{L}\!=\!j$ in $A$ of length $L$, and let the pivot weights on its L1L-1 pivot steps be $w_{r}:=W_{m_{r}m_{r+1}}\in[0,1]$ for $r=1,\dots,L-1$. Then
$(P_{K})_{ij}\;\geq\;\prod_{r=1}^{L-1}w_{r}.$
Proof. $P_{K}(i,j)$ is the probability of a union over all $\leq K$-hop paths; it is at least the probability of any particular path event. Under the model’s hop-1 semantics ($A$ gives certainty on the first hop), the path $p$ succeeds with probability $\prod_{r=1}^{L-1}w_{r}$. $\square$
Corollary 2 (uniform privileged pivots).
Under the conditions of Proposition 2, if every pivot step on $p$ satisfies $w_{r}\geq p$ for some $p\in[0,1]$, then
$(P_{K})_{ij}\;\geq\;p^{\,L-1}.$
Proof. By monotonicity of products, $\prod_{r=1}^{L-1}w_{r}\geq\prod_{r=1}^{L-1}p=p^{L-1}$; combine with Proposition 2. $\square$
5.7 Service-Class Lifting (Exact Land-then-Pivot Semantics)
Not all footholds are equal: landing via an interactive remote-service (e.g., RDP/SSH) typically enables immediate lateral movement, whereas landing via application protocols (e.g., SQL/HTTP) rarely does without an exploit. To encode this service memory exactly, we lift states from nodes to (node, class) pairs.
Classes and per-class adjacency.
@@ 5.7 Service-Class Lifting (Exact Land-then-Pivot Semantics)
$A\;=\;\bigvee_{c\in\mathcal{C}}A^{(c)},\quad\text{with }A_{ii}=0.$
Let $\pi_{\mathrm{pivot}}(c)\in[0,1]$ denote the pivot probability conferred by a foothold acquired via class $c$.
Lifted transition.
A lifted state is $(j,c)$ meaning “at node $j$ with foothold class cc.” From $(j,c)$ one can move to $(\ell,c^{\prime})$ if and only if $A^{(c^{\prime})}_{j\ell}=1$, and the step succeeds with probability $\pi_{\mathrm{pivot}}(c)$. This defines the class-conditioned transition matrix $T\in[0,1]^{(n|\mathcal{C}|)\times(n|\mathcal{C}|)}$ with blocks
$T_{(j,c)\to(\ell,c^{\prime})}\;=\;\pi_{\mathrm{pivot}}(c)\,\mathbf{1}\{A^{(c^{\prime})}_{j\ell}=1\}.$
Initialization (hop 1: pure reach).
From a start node $i$, the first hop ignores pivot and only records the landing class:
$P^{\mathrm{lift}}_{1}\big{(}i\to(j,c)\big{)}\;=\;A^{(c)}_{ij},\qquad P^{\mathrm{node}}_{1}(i\to i)=1.$
Iteration (hops $\geq 2$).
For $k\geq 1$, propagate in the lifted space using the same operators as in Section 3.3:
$P^{\mathrm{lift}}_{k+1}\;=\;P^{\mathrm{lift}}_{k}\;\lor_{p}\;\big{(}P^{\mathrm{lift}}_{k}\;\otimes\;T\big{)}.$
Aggregation back to nodes.
The $K$-hop node-to-node compromise probability is obtained by class-wise probabilistic union at the destination:
$(P_{K})_{i\ell}\;=\;\bigl{(}\;\lor_{p}\;\{\,P^{\mathrm{lift}}_{K}(i\to(\ell,c))\ :\ c\in\mathcal{C}\,\}\bigr{)}.$
All metrics (LMSK, BREK) continue to use $P_{K}$ exactly as defined here.
Compatibility with the base model.
If classes are collapsed by setting $W_{j\ell}=1-\prod_{c\in\mathcal{C}}\bigl{(}1-\pi_{\mathrm{pivot}}(c)\,\mathbf{1}\{A^{(c)}_{j\ell}=1\}\bigr{)}$, the lifted model reduces to the node-level model. Conversely, the lifted construction preserves hop-1 semantics and makes the dependence on the landing class explicit at every subsequent hop.
Theorems unchanged.
Proposition 1 (K-hop series upper bound) and Proposition 2 (single-path lower bound), and their corollaries, hold verbatim under this lifting. In Proposition 2 the per-step weights become the class-conditioned values $w_{r}=\pi_{\mathrm{pivot}}(c_{r})$ determined by the foothold class present at the source of step $r$. In Proposition 1, the path-sum bound is obtained by summing over class-typed paths; the resulting bound is still computable via ordinary matrix products by replacing $A$ with the block structure induced by $\{A^{(c)}\}$ and $W$ with $T$.

### v2

@@ 5.1 Lateral-Movement Susceptibility (LMSK)
Concept: This metric quantifies the overall ease of movement within the network. It answers the question: ”If we randomly select a directed pair $i\!\to\!j$ with $i\neq j$, what is the average probability that compromise propagates within $K$ hops?”
@@
Worked example (K=2 on the running graph).
From Fig. 7, the off-diagonal sum of $P_{2}$ is $4.17$; with $n=4$ there are n(n1)=12n(n{-}1)=12 off-diagonal entries, so
$\mathrm{LMS}_{2}\;=\;\frac{4.17}{12}\;\approx\;0.35.$
@@ 5.2 Blast-Radius Estimate (BREK)
Concept: This metric quantifies the expected scope of compromise. It answers: ”If the attacker starts at a uniformly random asset, what percentage of the network do we expect to be compromised?”
@@
$\mathrm{BRE}_{K}=\left(\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}(P_{K})_{ij}\right)\times 100\%$ (4)
@@
Worked example (K=2 on the running graph).
From Fig. 7, the total sum of $P_{2}$ is $8.17$ over $n^{2}{=}16$ entries, hence
$\mathrm{BRE}_{2}\;=\;\frac{8.17}{16}\times 100\%\;\approx\;51.06\%.$
@@ 6.  Formal Properties of the Model
Overview. We prove monotonicity, boundedness, and convergence; derive a $K$-hop series upper bound and a single-path product lower bound; recover boolean reachability in the unweighted limit; characterize convergence by the fixed-point stopping condition $P_{K+1}=P_{K}$ (equivalently, $P_{K}=P_{K}\oplus W_{K+1}$); under our dependence policy cycles add no mass, so stabilization occurs by $K\leq n{-}1$; and show all results persist under service-class lifting.
@@ 6.1 Monotonicity
This follows since $\oplus$ is monotone: with $W_{k+1}:=W_{k}\otimes W$, we have $P_{k+1}=P_{k}\oplus W_{k+1}$.
@@ 6.2 Boundedness
The values in the compromise probability matrix are always bounded within the interval $[0,1]$. The pivot potentials $\pi(s)$ are defined in $[0,1]$, and both the probabilistic composition ($\otimes$) and union ($\oplus$) operators are closed over this interval. This property guarantees the stability of the model, ensuring that the outputs are always valid probabilities.
6.3 Convergence (monotone bounded iteration)
Initialize $P_{1}:=I\oplus W$ and update layer by layer
$P_{k+1}\;:=\;P_{k}\;\oplus\;W_{k+1},\qquad W_{1}:=W,\;\;W_{t+1}:=W_{t}\otimes W\quad(t\geq 1).$
Each update is monotone entrywise (both $\oplus$ and $\otimes$ are monotone), and all values lie in $[0,1]$; hence the sequence $\{P_{k}\}$ is monotone and bounded and therefore converges. We denote the limit by
$P_{\infty}\;:=\;\lim_{K\to\infty}P_{K}.$
Stopping test. In practice we stop when
$P_{K+1}=P_{K}\quad\text{(within tolerance $\varepsilon$),}$
i.e., the (K+1)(K{+}1)-hop layer $W_{K+1}$ adds no new mass. Thus, at stabilization the matrix satisfies the fixed-point equation $P_{K+1}=P_{K}$ implies $W_{K+1}=0$ wherever $P_{K}<1$ (which holds off-diagonal in our model), since $P_{K+1}=P_{K}\oplus W_{K+1}$.
Dependence policy (edge sharing). We do not count as independent any walk that shares at least one edge with a shorter, already-accounted walk between the same endpoints; such a walk contributes no additional probability mass beyond the shorter walk.
Remark (cycles and truncation). Under the edge-sharing policy above, any walk that contains a cycle can be shortened by excising the cycle and therefore shares edges with a shorter walk; it adds no new independent mass. Enforcing strict simple-path semantics would require per-walk history and is exponential in the worst case [30]. In practice we use layered accumulation and stop early at a fixed point when $P_{K+1}=P_{K}$ (within tolerance), typically at $K\!\ll\!n$. As a hard upper bound, we truncate at K=n1K=n-1: any walk of length $\geq n$ necessarily repeats a vertex and thus contains a cycle [27, §1.3].
6.4 Connection to Algebraic Path Problems
We operate on $[0,1]$ with two primitives: a probabilistic union $\oplus$ and a probabilistic path composition $\otimes$. At the scalar level, $\oplus$ is the noisy-OR $a\oplus b=1-(1-a)(1-b)$ and $\otimes$ is an abstract operator for sequential steps (under conditional independence one may take $a\otimes b:=ab$). The matrix lifts are
$(X\oplus Y)_{ij}=X_{ij}\oplus Y_{ij},\qquad(X\otimes Y)_{ij}=\bigoplus_{m=1}^{n}\bigl(X_{im}\otimes Y_{mj}\bigr).$
These operators are monotone on $0,1]^{n\times n}$; $\oplus$ is commutative and associative (non-idempotent), and in our setting the super-distributive bounds hold
$(X\otimes Y)\oplus(X\otimes Z)\;\geq\;X\otimes(Y\oplus Z),\qquad(Y\otimes X)\oplus(Z\otimes X)\;\geq\;(Y\oplus Z)\otimes X.$
The layered update $P_{k+1}=P_{k}\oplus W_{k+1}$ thus computes an algebraic-path closure by monotone iteration; at stabilization the result $P$ satisfies the fixed-point equation $P=P\oplus(P\otimes W)$, analogous to algebraic path problems via least fixed-point iteration on a monotone operator [21, 22, 13, 17].
@@ 6.5 Refinement of Binary Reachability
Let $R_{K}$ be the final reachability matrix from the unweighted (Boolean) method, and $P_{K}$ be the final matrix from our probabilistic method. The following soundness relationship holds:
$(P_{K})_{ij}>0\;\Rightarrow\;(R_{K})_{ij}=1$
@@
6.6 Bounds for $K$-hop Lateral Propagation
Upper bound by Boolean reachability (trivial).
Let $R_{K}$ be the Boolean $K$-hop reachability matrix. Then for all $i\neq j$,
$0\;\leq\;(P_{K})_{ij}\;\leq\;(R_{K})_{ij}\;\in\{0,1\},\qquad\text{and}\qquad(P_{K})_{ii}=1.$
Reason. If $(P_{K})_{ij}>0$ then some $\leq K$ walk exists from $i$ to $j$, hence $(R_{K})_{ij}=1$. The diagonal equality follows from $P_{1}=I\oplus W$.
@@ 6.6 Bounds for $K$-hop Lateral Propagation
Fix $K\geq L\geq 1$. Suppose there exists a directed path (no repeated vertices)
$p:\ i=v_{0}\to v_{1}\to\cdots\to v_{L-1}\to v_{L}=j,$
with step-successes $w_{r}:=W_{v_{r-1}v_{r}}\in[0,1]$ for $r=1,\dots,L$. Then
$(P_{K})_{ij}\;\geq\;w_{1}\,\otimes\,w_{2}\,\otimes\,\cdots\,\otimes\,w_{L},$
and under the scalar instantiation $a\otimes b:=ab$ this equals $\prod_{r=1}^{L}w_{r}$.
Proof.
  • By construction, $W_{1}=W$ and $W_{t+1}=W_{t}\otimes W$ expand all length-(t+1)(t{+}1) walks; hence the specific path $p$ contributes at least $w_{1}\otimes\cdots\otimes w_{L}$ to $(W_{L})_{ij}$.
  • Since $P_{K}=I\oplus W_{1}\oplus\cdots\oplus W_{K}$ and $L\leq K$, monotonicity of $\oplus$ implies
    $(P_{K})_{ij}\;\geq\;(W_{L})_{ij}\;\geq\;w_{1}\otimes\cdots\otimes w_{L}.$
$\square$
6.7 Service-Class Lifting via Node Splitting (Land–then–Pivot)
The core idea of our advanced model is to distinguish how a foothold is landed (the service used to gain initial access to a node) from the service used for the next hop. These “land–then–pivot” semantics reflect a real-world observation: an attacker’s power to pivot from a server depends on how they landed on it in the first place.
For example, if an attacker lands via a high-privilege service like SSH, the next hop succeeds with a high probability $\pi(\mathrm{SSH})$ as attacker needs sufficient privilege to make the next hop. Conversely, if the attacker lands via a low-privilege service like MSSQL, the next hop succeeds with a much lower probability $\pi(\mathrm{MSSQL})$. Thus, the hop success probability is a function of the landing service, not the outgoing service. This “memory” of the landing service is key to a more precise model.
We retain a single adjacency $A\in\{0,1\}^{n\times n}$ on nodes $V$, and attach to each node $v\in V$ a small set of services $S(v)$ (e.g., $S(i)=\{\mathrm{SSH},\mathrm{MSSQL}\}$, $S(j)=\{\mathrm{SSH}\}$, …). Each service $s$ has a pivot rate $\pi(s)\in[0,1]$ that reflects the power an attacker has because they landed via $s$.
@@ 6.7 Service-Class Lifting via Node Splitting (Land–then–Pivot)
$A=\begin{bmatrix}0&1&1&0\\ 0&0&1&1\\ 0&0&0&1\\ 0&0&0&0\end{bmatrix}$
Figure 8: Directed graph and its adjacency matrix AA.
$S(i){=}\{\mathrm{SSH},\mathrm{MSSQL}\},\quad S(j){=}\{\mathrm{SSH}\},\quad S(k){=}\{\mathrm{MSSQL}\},\quad S(\ell){=}\{\mathrm{SSH}\},$
and rates
$\pi(\mathrm{SSH}){=}0.8,\quad\pi(\mathrm{MSSQL}){=}0.3.$
Lifted state space and transition.
The lifted vertices are $(v,s)$ with $s\in S(v)$. We define a single lifted transition matrix
$W^{\uparrow}_{(u,s_{\mathrm{in}})\to(v,s_{\mathrm{out}})}\;=\;\pi(s_{\mathrm{in}})\;\mathbf{1}\{A_{uv}=1,\ s_{\mathrm{out}}\in S(v)\}.$
This ties the hop success to the landing service carried by the source state.
In lifted order $(i,\mathrm{SSH}),(i,\mathrm{MSSQL}),(j,\mathrm{SSH}),(k,\mathrm{MSSQL}),(\ell,\mathrm{SSH})$,
$(i,\mathrm{SSH})$$(i,\mathrm{MSSQL})$$(j,\mathrm{SSH})$$(k,\mathrm{MSSQL})$$(\ell,\mathrm{SSH})$0.80.80.80.80.30.30.30.30.80.80.80.80.30.3
$W^{\uparrow}=\begin{bmatrix}0.0&0.0&0.8&0.8&0.0\\ 0.0&0.0&0.3&0.3&0.0\\ 0.0&0.0&0.0&0.8&0.8\\ 0.0&0.0&0.0&0.0&0.3\\ 0.0&0.0&0.0&0.0&0.0\end{bmatrix}$
Figure 9: Lifted state graph $V^{\uparrow}=\{(v,s):s\in S(v)\}$ with transition weights equal to the pivot rate of the landing service.
Initialization (lateral-only) and iteration in lifted space.
Let $I^{\uparrow}$ be the identity on lifted states and $W^{\uparrow}$ the one-hop lifted operator. Define
$W^{\uparrow}_{1}:=W^{\uparrow},\qquad W^{\uparrow}_{k+1}:=W^{\uparrow}_{k}\otimes W^{\uparrow}\quad(k\geq 1).$
Accumulate the $K$-hop series using the usual operators:
$P^{\uparrow}_{1}:=I^{\uparrow}\oplus W^{\uparrow}_{1},\qquad P^{\uparrow}_{k+1}:=P^{\uparrow}_{k}\oplus W^{\uparrow}_{k+1}\quad(k\geq 1).$
$(i,\mathrm{SSH})$$(i,\mathrm{MSSQL})$$(j,\mathrm{SSH})$$(k,\mathrm{MSSQL})$$(\ell,\mathrm{SSH})$11111111110.80.80.80.80.30.30.30.30.80.80.80.80.30.3
$P^{\uparrow}_{1}\;=\;I^{\uparrow}\oplus W^{\uparrow}\;=\;\begin{bmatrix}1.0&0.0&0.8&0.8&0.0\\ 0.0&1.0&0.3&0.3&0.0\\ 0.0&0.0&1.0&0.8&0.8\\ 0.0&0.0&0.0&1.0&0.3\\ 0.0&0.0&0.0&0.0&1.0\end{bmatrix}$
Figure 10: Lifted one-hop reach $P^{\uparrow}_{1}$ (identity + one-hop).
Collapse to nodes (single, consistent rule).
After computing reachability in the lifted space, we return to node–to–node probabilities by a single noisy–OR over both the source and destination service labels:
$(P_{K})_{uv}\;=\;\bigoplus_{s\in S(u)}\;\bigoplus_{t\in S(v)}\;P^{\uparrow}_{K}\!\big[(u,s)\to(v,t)\big].$
We always apply this double union, for all $K$, and we perform it once at the end (i.e., no intermediate collapses). This corresponds to the modeling choice that an attacker can leverage any foothold service at the source node and, on a hop, can attempt all destination interfaces; independence across attempts is captured by the noisy–OR operator $\oplus$. Since $P^{\uparrow}_{K}$ includes $I^{\uparrow}$, node diagonals of $P_{K}$ are handled automatically.
[…]
$P_{1}\;=\;\begin{bmatrix}1.00&0.86&0.86&0.00\\ 0.00&1.00&0.80&0.80\\ 0.00&0.00&1.00&0.30\\ 0.00&0.00&0.00&1.00\end{bmatrix}$
Figure 11: Node-level one-hop reach $P_{1}$ (identity + one-hop).
$(i,\mathrm{SSH})$$(i,\mathrm{MSSQL})$$(j,\mathrm{SSH})$$(k,\mathrm{MSSQL})$$(\ell,\mathrm{SSH})$0.640.640.730.730.240.240.310.310.240.24
$W^{\uparrow}_{2}\;=\;\begin{bmatrix}0.00&0.00&0.00&0.64&0.73\\ 0.00&0.00&0.00&0.24&0.31\\ 0.00&0.00&0.00&0.00&0.24\\ 0.00&0.00&0.00&0.00&0.00\\ 0.00&0.00&0.00&0.00&0.00\end{bmatrix}$
Figure 12: Lifted exactly-two-hop transition $W^{\uparrow}_{2}=W^{\uparrow}\otimes W^{\uparrow}$.
$(i,\mathrm{SSH})$$(i,\mathrm{MSSQL})$$(j,\mathrm{SSH})$$(k,\mathrm{MSSQL})$$(\ell,\mathrm{SSH})$11111111110.80.80.930.930.730.730.30.30.470.470.310.310.80.80.850.850.30.3
$P^{\uparrow}_{2}\;=\;P^{\uparrow}_{1}\oplus W^{\uparrow}_{2}\;=\;\begin{bmatrix}1.00&0.00&0.80&0.93&0.73\\ 0.00&1.00&0.30&0.47&0.31\\ 0.00&0.00&1.00&0.80&0.85\\ 0.00&0.00&0.00&1.00&0.30\\ 0.00&0.00&0.00&0.00&1.00\end{bmatrix}$
Figure 13: Lifted reach $P^{\uparrow}_{2}$ (identity + one-hop + two-hop union).
[…]
$P_{2}\;=\;\begin{bmatrix}1.00&0.86&0.96&0.81\\ 0.00&1.00&0.80&0.85\\ 0.00&0.00&1.00&0.30\\ 0.00&0.00&0.00&1.00\end{bmatrix}$
Figure 14: Node-level reach $P_{2}$ (identity + union of all 1–2 hop paths via lifted space).
What this shows.
We use a single model: node splitting by service with land–then–pivot semantics. Hop success depends on the landing service carried by the source state, so all compositions are done in the lifted space and we collapse back to nodes only once via a noisy–OR over both source and destination service labels. This preserves landing–service memory across hops; any intermediate collapse to nodes would lose that memory and undercount reach. The resulting $P_{K}$ feeds directly into $\mathrm{LMS}_{K}$ and $\mathrm{BRE}_{K}$.